Are K=10, C=1 valid witnesses to show that f(x)=10 is O(x)?
Answer #1
f(n) = O(g(x)) means there are positive constants c and k, such that 0 ≤ f(x) ≤ cg(x) for all x ≥ x0

10 = O(x)
=>  10 <= cx
Let's assume c = 1
=>  10 <= cx
=>  10 <= 1*x
=>  10 <= x
it's true for all x >= 10

so, 10 = O(x) given c = 1 and k = 10